Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

NIght Report 2015-02-15

Observers : Laurent Le Guillou, Ayan Mitra

Support team : Nicolas Chotard, Pierre Francois Leget 

TAT : Joe Anderson, Nancy Elias-Rosa, Michel Dennefeld 


Clear sky, low wind, fluctuating humidity (but below 55%).

Seeing at 0.7' at the beginning of the night.


(times are in UT)


CALIBRATIONS :

 - Evening : EFOSC_bias, EFOSC_FlatSpec, EFOSC_SpecArc

 - Morning : EFOSC_SkyFlat: grz


SCIENCE OBSERVATIONS :


--- SOFI ---


 - 00:03  ASASSN-14ha photometry with J 

                   

 - 00:50  LTT3218 flux standard BG RG 


 - 00:51  seeing ~ 0.8" , humidity 55%, temperature 15.6°C


 - 01:03 HIP020378 telluric standard BG RG 

 

 - 01:09 ASASSN-14ha  spectroscopy RG BG 

            aborted during BG no more signal 


 - 03:40 HIP020378 telluric standard BG RG 


 - 03:48 LTT3218 flux standard BG RG


 - 04:04 HIPO53172 telluric standard BG RG


 - 04:10 PSNJ10544216  BG RG 


 - 05:57 seeing ~ 1.1, humidity 28%, temperature 16°C


 - 06:43 HIPO53172 telluric standard BG RG 


 - 06:56 PSNJ10544216 photometry JHK 


--- EFOSC ---


 - 07:26 LSQ15kp follow up photometry 

         

 - 07:56 LSQ15kp spectroscopy g11+g16 1800s each.


 - 09:00 LTT3864 flux standard g11+g13+g16               

          


 


Comments